Solve the system by substitution.
4x - 4y = -16
y = 5x - 8

Answer by checkley77(12844)

4x - 4y = -16
y = 5x - 8
4X-4(5X-8)=-16
4X-20X+32=-16
-16X=-16-32
-16X=-48
X=-48/-16
X=3 ANSWER.
Y=5*3-8
Y=15-8
Y=7 ANSWER.
PROOF:
4*3-4*7=-16
12-28=-16
-16=-16